THE PERCENTAGE BY VOLUME OF PROPANE IN A MIXTURE OF PROPANE,METHANE AND CO IS 36.5 CALCULATE THE VOLUME OF CARBON DIOXIDE PRODUCED WHEN 100 ML OF THE MIXTURE IS BURNT IN EXCESS OF OXYGEN?

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi,

propane molecular formula is C3H8 which gives 3 ml of CO2 per 1 mole

But the CO and CH4 give one mole each per one mole.

In the 100 ml mixture the CO and CH4 are (100-36.5)=63.5 ml which will give only 63.5 ml of CO2

And propane gives 3×(36.5)ml of CO2.

Total CO2 produced = 109.5 + 63.5 = 173 ml.

chandra sekhar

askiitian.expert